OpenStudy (anonymous):

Find the equation of the circle that is tangent to the line whose equation is given by x + y = 2 and has its center at (3,5).

OpenStudy (anonymous):

This is another version of finding the perpendicular to a line given a point and a line. Then the circle will be x^2+y^2 = a^2 where a is the radius (distance)

OpenStudy (anonymous):

real job here is to find the distance between the point and the line yes?

OpenStudy (anonymous):

@estudier i think the equation will be \[(x-3)^2+(y-5)^2=r^2\] we just need r

OpenStudy (anonymous):

(x-3)+(y-5)=20

OpenStudy (anonymous):

Yes, just a translation..

OpenStudy (anonymous):

forgot to square the terms my bad

OpenStudy (anonymous):

i am trying to think of a snap way to get the distance without dropping a perpendicular to the line from the point and then computing. how did you get \[\sqrt{20}\]?

OpenStudy (anonymous):

Midpoint of the line is the point where it would be tangential due to the slope being 1, so find the midpoint of the line (from x to y axis, I had (1,1), you should verify this, been up for 24 hours :S) and then use the distance formula from the center of the circle to the midpoint of the line. I may have gone wrong somewhere, but that was my methodology.

OpenStudy (anonymous):

I agree though, your method would be better :). Find the perpendicular line first, then you can find a point on the line and the circle, follow up with the distance formula, and wholla your answer :).

OpenStudy (anonymous):

i can do it as follows, but i thought there might be a snap way. a) the slope of the line x+y=2 is -1 so the slope of the line giving the distance from (3,5) to x +y=1 will be -1 2) the point on the line will be (x, 2-x) 3) so we find the piont via \[\frac{5-(2-x)}{3-x}=1\] \[\frac{3+x}{3-x}=1\] \[x=0\] and so the point on the line closest to (3, 5) is (0,2)

OpenStudy (anonymous):

but if i do this i don't get 20 for the square of the distance, get 18

OpenStudy (anonymous):

U could translate the circle and line to (0,0) -> Pythagorus.

OpenStudy (anonymous):

maybe i messed up

OpenStudy (anonymous):

@estudier i think easier just to find square of distance because we already have center. all we need is r^2 to write equation

OpenStudy (anonymous):

No, I agree with you, I guess I made a bad assumption, should've actually solved it the long way :).

OpenStudy (anonymous):

oh ok. so answer is \[(x-3)^2+(y-5)^2=18\]

OpenStudy (anonymous):

@kris make any sense?

OpenStudy (anonymous):

i got. (x - h)2 + (y - k)2 = r2 (x - 3)2 + (y - 5)2 = (3sqrt(2))r2 (x - 3)2 + (y - 5)2 = 18

OpenStudy (anonymous):

Just for the sake of it, take the circle at (0,0), x^2 +y^2 = r^2 and we want r. Translate the line -> x+3+y+5 =2-> x + y =-6 After this, you have a right triangle at the origin bounded by the line with sides of 6 and r is the altitude so r^2 = 6^2 - sqrt(72/2)^2 = 18 and r = sqrt 18